OpenStudy (sissyedgar):

What is the solution to the linear inequality −1/4x≤8? Enter your answer in the box. l______l ♥ Thank you! ♥

alones (alones):

Multiply both sides by \[\frac{ 4 }{ (-1) }\]

alones (alones):

\(\frac{ 4 }{ -1}*\frac{ -1 }{ 4 }x\)≤\(\frac{ 4 }{ -1}*(8)\)

alones (alones):

\(\ x≥−32\) Should be your answer.

OpenStudy (sissyedgar):

So